OpenStudy (anonymous):

Guys, this is the question I have in Math: How is an inequality different from an equation? Give a real-world scenario in which you would write an inequality rather than an equation. So, I'm not very good at math, and if you could help, that would be great. :D

OpenStudy (igreen):

Inequality signs are \(<, \le, >, \ge\). \(<\) is the less than sign. It says that the value to the left of it is less than the value to the right of it. Example: \(3 < 5\). \(\le\) is the less than or equal to sign. It says the value to the left of it is less than or the same as the value to the right of it. Example: \(x \le 6\) This means that \(x\) has to be either 6 or a number less than 6. \(>\) is the greater than sign. It says that the value to the left of it is greater than the value to the right of it. Example: \(9 > 2\) \(\ge\) is the greater than or equal to sign. It says the value to the left of it is greater than or the same as the value to the right of it. Example: \(x \ge 4\) This means that \(x\) has to be either 4 or a number greater than 4. \(=\) is the equal sign. It says that 2 values are the same.

OpenStudy (igreen):

In a world-case scenario, your customer asks for more than 2 boxes of your product. You can't use the equal sign, because he says he wants more than 2 boxes, so you must use the inequality sign \(>\).

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @LyssH